write your answer as an ordered pair.
y=x+3 y=2x-1

Answer by Cromlix(4381)

y = x + 3.....1
y = 2x - 1....2
Substitute x + 3 for y in equation(2)
x + 3 = 2x - 1
x - 2x = - 1 - 3
-x = -4
x = 4
Substitute x = 4 in equation (1)
y = 4 + 3
y = 7
{4,7}
Hope this helps.
:-)